Day To Day
Our Beeyard Journal
As well as being very hands-on with our bees, we continually learn from other beekeepers, attend an annual 'bee school', and stay up to date on beekeeper news through our two club memberships.
See what we're doing (read below).
<--- our humble beeyard
Questions? Want To Visit?
Support our live pollinator habitat
[ click any photo to enlarge ]
Beekeeper's Diary 2021
> bee husbandry is 24/7
00Feb21 : next!
20Feb21 : Seven days of snow and temps of 11°-32°! Yesterday a warming started and today was 55°. The all hives has been busy taking 'cleansing' flights and cleaning house with the Sterling hive taking a particularly bad hit. They all have been stressed so I will add top-feeding support to each once temps are consistently 55°+.
14Feb21 : 27°and winter gray with a Valentines ice storm. Tarped hives on 3 sides last night before ice storm moved in. With our small beeyard, we treat them like any livestock, meaning they can be cold or wet but avoid cold AND wet.
We also don't want their reduced entrances, the ONLY source of air flow, to ice over. On days over 32°, I have been feeding but 5-6 days below 30° will be a severe test. Fearing the worst, I have ordered a 2021 nuc.
24Jan21 : feeding hives; cleaning and repairing gear; restringing frames for just wax & honeycomb harvests.
11Jan21 : feeding today as it is sunny and about 50°! Just a quick lift of lids and dropping in some bee food while we see how each hive is faring. Eleanor had a huddle of bees in their winter board (kinda unusual but ok); Sterling still had bee food (assuming their stores are sustaining them ok as they had lots of entrance activity; and none had any condensation issues. We have been purchasing, repairing, and repairing gear for the Spring up-tick starting next month!
So far, it looks like we have all hives surviving the Winter (zero winter kill)! So, dunno if we need to order a spare nuc > the longer we wait to order, the harder to find a source.
9Dec20 : time to install those winter boards (to help control hive-killer condensation/moisture)!
6Dec20 : oxalic acid treatment (for varroa mites) for all hives; Sterling has lots of entrance activity while others seem low key. Put up the winter wind barrier.
21Nov20 : light feeding on warm days.
14Oct20 : warm, bright/clear and 80°
We are heading to the beeyard to remove the strips of Formic Acid which have been installed for 2 weeks now. We also have chores to clean, reorganize and refurbish the beeyard for winter and 2021. This means removing the shims, inspecting bees, treating for pests, checking honey stores, painting supers, and prepping empty supers.
2Oct20 : warm, overcast and 79°
Time for 2nd mite treatment of the year. We are using a sticky cheese-like strip of gel called Formic Acid which is placed on top of the brood super/box.
11Sep20 : warm again, bright/clear and 90° at 1pm.
We enjoyed cool evenings over Labor Day weekend but summer is back for one last blaze. Good to see a bit of Fall flow going on that we didn't have last year. All hives are healthy and we plan to take all 3 into winter (will not merge hives). Not feeding now but letting girls clean up some frames and wax.
18Aug20 : tad humid, breath of breeze, clear, sunny 74° at 8am
3 days ago gave Eleanor and Sterling entrance feeders to hold them until today when I replenish the feeding supers on all. I totally believe that the feeding supers help draw heat out of the hive but there is still serious bearding every afternoon.
CANNOT SAY enuff good things about using a fume board (right) instead of a smoker! OMG, in 2 minutes bees have moved down into super below. This is like a tall inside cover > you just spray it with BeeBGone, cover the super, and the girls move out of your way, yay! I also feel like it lowers the risk of me getting in the queen's way while I inspect frames - whew!
This August is not nearly as hot and dry as last year, so feeding schedule and amount is just to augment and support hives, not sustain. All hives seem to be doing very well. Sterling remains the smallest hive and doesn't seem to be growing into their top super so it was removed to reduce risk of pests.
20Jul20 : Sunny but drippy humid and already 82° at 9am
Apriguard treatment #2 for our 3 hives. Every hive looks strong with no pest issues and has at least 1 heavy super for winter. All used at least 1 full quart of their internal syrup feeder since the 11th. Internal feeding seems supremely better than external, and I like that the super with the mason jars can also help some of the excessive heat from lower down escape. Sterling remains the questionable -- strong but not exceedingly strong like the other two. Will add entrance reducers tomorrow to help reduce robbing risk.
11Jul20 : Sunny, humid, 94° | brief afternoon showers, a tad warmer than last year
Busy day > final honey harvest on Eleanor, Cloud9, and Sterling -- hope to get at least 24 fully capped frames. (Thugs colony did not accept new queen and is officially vacated, so will close this hive -- back to 3 hives but these are the strongest ever!) So, today we will break down the 3 hives and apply 1st mite treatment, Apiguard, then restack like this:
deep super on inner cover holding 2 feeding jars (2:1 syrup*)
shim holding Apiguard mite treatment
deep brood super
2nd Apiguard treatment in 10-14 days.
the shim which leaves a 1" space for bees to come in contact with treatment gel).
Cloud9 and Sterling, being new hives, have never had a mite treatment -- 2nd Apiguard treatment in 10-14 days.
*4 cups water = 2# to 1# sugar or 24 cups water and 12# sugar to make the 6 quarts we need.
The Thugs hive has been abandoned/failed and moths moved in so we also will clean that up.
3Jul20 : Harvested 4 frames from Cloud9 weighing 5# each!
12Jun20 : Cloud9 was filling supers fast at last inspection so I know they need room but 5 boxes is enough as a 6th will be for internal feeding soon, so time to harvest a bit. And what a great harvest of just 9 frames! (not even a whole super)! This pix is about 70% of what we got today! Cloud9 is a very healthy and robust hive; textbook perfect is every respect.
----> COVOD has bee clubs meeting virtually or requiring masks with temp checks
26May20 : transporting 2 new queens for Tink/Thug hives (video below); maybe a new queen will have nicer brood forcing me to rename Thugs; all Eleanor offspring are very docile bees; added 4th box to Cloud 9 ('Eleanor' swarm of 29Mar).
25May20 : Tink has had 'Eleanor' split added for over 3 weeks but population is way down (only a few older brood), so guessing hive is queenless - same scenario for Thugs.
25May20 : got 2 MORE frames of honey from Eleanor
9May20 : got first 2020 honey, 2 frames of honey from Eleanor
1May20 : combined 'Eleanor' split to strengthen Tink hive
24Apr20 : 8-day old Tink swarm inspection (video below)
15Apr20 : new swarm ("Tink") is fine and highly productive in 1 week! Ordered supers to ensure we can grow FIVE hives and test internal feeding this summer (use deep supers for mason jar sugar-water feeders). Beeyard now has 4 colonies!
----> COVOD has bee clubs cancelled
29Mar20 : caught huge swarm ("Cloud 9") from Eleanor which caught ME in its swarm cloud and landed shoulder-height in the tree right beside me! It took FIVE 2-gal buckets to grab all the bees.
27Mar20 : Eleanor is huge colony and overproducing drone so split hive
10Feb20 : both colonies totally out of fondant; Tz'U Hsi (aka thugs) has massive condensation under cover; wiped off; both colonies show strong foraging but on different days
23Jan20 : 2nd oxalic treatment for Tz'U and Eleanor hives & fondant
16Jan20 : 1st oxalic treatment for Tz'U Hsi and Eleanor hives & fondant
30Oct19 : ends Formic acid treatment cycle
23Oct19 : new queen for Eleanor hive; winterize beeyard
22Oct19 : reintroduced Zena back into original Tz'U Hsi hive
18Oct19 : Formic acid treatment for Tz'U Hsi and Eleanor hives
29Aug19 : ended Apiguard treatment cycle
3Aug19 : 2nd Apiguard treatment for Tz'U Hsi and Eleanor hives
21Jul19 : 1st Apiguard treatment of the year
31May19 : caught Tz'U Hsi swarm creating Zena
27Apr19 : got another 5-frame nuc from Bethpage area.
10Apr19 : got 5-frame nuc from Sterling.
1Jan19 : hive that was robbed last fall did not survive.